Find an equation for the ellipse. Graph the equation.
Center at (5,4); focus at (12,4); contains the point (5,9)

Answer by Boreal(15235)

It will have (x-5)^2 and (y-4)^2 as its center.
focus is along the x-axis.
(5,9) is the point at one end of the semi-minor axis
a^2-c^2=b^2
74-49=25
a=sqrt(74)
Major axis from (13.6,4) to (-3.6,4)
Minor axis from (5,-1) to (5,9)
Foci at (12,4), (-2,4)
[(x-5)^2/74] +[ (y-4)^2/25] =1